OpenStudy (anonymous):

Can somebody please explain to me how to do this? A plane takes a trip at 65 mi/h. A plane traveling 30 mi/h makes the same trip in 3 fewer hours. Write and solve an equation to find the distance of the trip.

hartnn (hartnn):

you basically need to use just one equation : \(speed \times time=distance\)

hartnn (hartnn):

for first plane, whats your equation, can you form it ?

hartnn (hartnn):

let 1st plane take 't' hours, so 2nd plane will take t-3 hours, right ? so, the 2 equations you get are : 65 *t = d 30 (t-3) =d can you solve these 2 equations for d ?

OpenStudy (anonymous):

how can I solve for d if I don't know what t is?

hartnn (hartnn):

you have 2 equations and 2 unknowns

hartnn (hartnn):

65 *t = d 30 (t-3) =d so, put d= 65t from 1st equation into 2nd 30(t-3) = 65d find 't' from here, then d is just 65t

OpenStudy (anonymous):

OHH thank you!!

hartnn (hartnn):

welcome ^_^